# Action Spotting and Precise Event Detection in Sports: Datasets, Methods, and Challenges

### v1

@@ 3.1 Temporal Action Localization
Table 2: Summary of methods for Action Spotting (AS) and Precise Event Spotting (PES). Performance is grouped under the Test and Challenge sets. All results are reported on the SoccerNet Action Spotting dataset, where italicized entries indicate results from SoccerNet-v1 [14], and plain text entries are from SoccerNet-v2 [29]. Bold numbers indicate the highest scores in each column. A ✓in the generalizability column means the method was tested on other sports beyond soccer.
Method Year Technique Modality Test Set Challenge Set Generalizability
@@
Giancola et al. [14] 2018 Temporal Pooling Visual 31.37 30.74
Rongved et al. [30] 2020 3D Convolution Visual 56.3
Vats et al. [31] 2020 Multi-Temporal Scales Visual 60.1
CALF [32] 2020 Context-aware Loss Visual 41.61 42.22
Vanderplaetsen et al. [33] 2020 Multi-Modal Fusion Visual + Audio 39.90
NetVLAD++ [34] 2021 Temporally Aware Pooling Visual 53.40 52.54
RMS-Net [35] 2021 Regression & Masking Visual 63.49 60.92
Zhou [36] 2021 Transfer Learning Visual 47.05 74.77 49.56 74.84
E2E-Spot [15] 2022 GRU + Gated Shift Module Visual 66.73 73.62
Shi et al. [37] 2022 Multiple Scene Encoders Visual 55.20
STE [38] 2022 1D Convolution Visual 58.29 71.58 58.71 70.49
SpotFormer [39] 2022 Multi Feature Extractors Visual 60.90 81.50
Soares et al. [40] 2022 Dense Detection Anchor Visual 65.07 78.59 68.33 78.06
Zhu et al. [41] 2022 Multi-Scale Vision Transformer Visual 52.04 60.86
ASTRA [42] 2023 Transformer Encoder-Decoder Visual + Audio 70.10 79.21
T-DEED [43] 2024 Scalable-Granularity Perception Visual
COMEDIAN [44] 2024 Self-Supervised Learning & Knowledge Distillation Visual 73.10 68.38 73.98
Tran et al. [45] 2024 Vision-Language Object Detector Visual + Text 62.49 73.98 69.38 76.15
3.2 Action Spotting & Precise Event Spotting
In contrast to TAL, which focuses on predicting temporal intervals for actions, AS and PES have gained significant attention in the sports analytics domain. AS and PES aim to identify actions or events using a single key frame, making them particularly suited for rapidly occurring events while also reducing annotation effort [43, 46]. This contrasts with traditional temporal video understanding, where actions are defined by explicit start and end times. In sports scenarios, these boundaries often become ambiguous—for example, clearly defining the precise start and end of a goal event can be challenging.
To address the ambiguity in defining temporal boundaries in sports actions, the SoccerNet dataset introduced AS as a new task specifically tailored to the sports domain [14], as illustrated in Figure 4. Alongside the dataset, S. Giancola et al. also proposed the first baseline method for AS, which involved evaluating various temporal pooling strategies on pre-extracted features from video chunks. Building on this initial work, they later introduced NetVLAD++ [34], a temporally-aware pooling module inspired by NetVLAD [47]. While standard NetVLAD aggregates frame-level features without considering temporal order—treating frames as an unordered set—NetVLAD++ explicitly incorporates temporal structure. This is particularly important in AS scenarios, where the sequence of frames before and after an event carries significant contextual information. By modeling both past and future contexts, NetVLAD++ achieved improved performance in AS.
A summary of recent methods developed for AS and PES, including their key techniques, modalities, and evaluation results on the SoccerNet benchmark, is provided in Table 2.
O. Rongved et al. [30] trained a 3D ResNet model, inspired by the architecture proposed in [48], and applied it directly to raw video frames using a 5-second sliding window approach. While 3D CNNs are effective at capturing spatiotemporal features, they are also known for their high computational cost, which can hinder real-time deployment and scalability. To address this limitation, Darwish et al. introduced the Spatio-Temporal Encoder (STE) [38], which relies solely on 1D convolutions and fully connected layers. This design significantly reduces computational complexity while maintaining strong performance, demonstrating that efficient architectures can still effectively model spatiotemporal dynamics without the overhead of 3D convolutions.
K. Vats et al. [31] proposed a multi-tower CNN architecture designed to process video information at multiple temporal scales, addressing the uncertainty in AS. In parallel, Cioppa et al. [32] introduced a context-aware loss function that explicitly models the temporal context surrounding actions. Their method takes a different approach by predicting multiple action spots within each video chunk, achieved by regressing several temporal offsets per chunk to capture the potential positions of events.
A significant performance improvement was achieved by Zhou et al. [36], who demonstrated that task-specific feature representations tailored to soccer broadcast videos can substantially enhance the effectiveness of AS models. They fine-tuned several pre-trained action recognition networks on soccer-specific data to extract high-level semantic features that better reflect the visual patterns and context of the sport. These features were then fed into Transformer-based encoding layers serving as the prediction head. Their approach achieved SOTA at the time, highlighting the importance of leveraging domain-adapted features for improving spotting accuracy in soccer video analysis.
Building on detection principles and self-supervised learning, Tomei et al. [35] presented RMS-Net, a model that combines regression and masking strategies. Inspired by object detection pipelines and pretraining techniques, RMS-Net enables the network to learn robust representations for spotting actions with improved precision.
J. Hong et al. later extended AS to PES by setting a stricter tolerance window [15]. Their proposed E2E-Spot model begins by extracting local spatial-temporal features using RegNet-Y [49] enhanced with Gate Shift Modules [50]. These features are passed through a single-layer bidirectional GRU [51], which outputs per-frame event occurrence probabilities.
In parallel, methods originally developed for TAL have shown strong performance when adapted to AS tasks. For example, J. Soares et al. proposed a detection approach based on short temporal anchors (e.g., 0.5–1.0s) where each anchor is a pair formed by a time instant and action class, predicting confidence scores and fine-grained temporal displacements for each anchor [40]. Their framework, inspired by object detection, employs both a 1D U-Net [52] and a Transformer-based encoder, with the Transformer outperforming by effectively modeling both temporal and contextual information [53].
Incorporating multi-modal input, especially audio, has also proven beneficial as audio data can contain many critical cues such as the sound when the player strikes the ball. Vanderplaetsen et al. [33] explored different fusion strategies and found that late fusion—integrating audio and visual features just before the fully connected layer—yielded the best performance.
With the rise of Transformers in computer vision [54], their application has also extended to the PES task. SpotFormer [39] adopts this approach by leveraging multiple feature extractors—including Transformer-based models (VideoMAE [55], SwinTransfrmer [56])—to extract rich spatial and temporal features from soccer videos. These features are then fed into a Transformer-based spotting network to perform precise event spotting. This multi-extractor strategy achieved strong performance, demonstrating the effectiveness of Transformer architectures in learning from diverse feature inputs. However, the use of multiple feature extractors significantly increases computational cost and inference time, making it less practical for real-time or resource-constrained applications.
To address temporal modeling more efficiently, Zhu et al. [41] proposed the use of a multi-scale Vision Transformer for feature extraction, followed by a sliding window mechanism to improve temporal understanding. These enhanced temporal features are then processed using NetVLAD++, a pooling strategy designed to aggregate local features into a robust global representation, producing the final event spotting predictions. This approach strikes a balance between temporal context modeling and computational efficiency, offering an alternative to the heavier multi-extractor pipelines.
In a similar direction, Shi et al. [37] introduced a multi-scene encoding strategy to address the variability in action durations. First, video features are extracted as a chunk and then divided into multiple subsets, each of which is processed by separate Transformer encoders. By varying chunk sizes, their model captures actions of different temporal lengths more effectively, demonstrating a flexible and adaptive approach to temporal modeling in event spotting tasks.
Building on the success of transformer and audio fusion, Xarles et al. proposed ASTRA (Action Spotting TRAnsformer) [42], a Transformer-based encoder-decoder model designed for soccer videos. ASTRA processes both visual and audio embeddings using learnable queries in the decoder, demonstrating significant performance gains on SoccerNet by leveraging multi-modal information with transformers.
Tran et al. [45] highlight a key limitation in traditional backbone models: they tend to rely heavily on global feature extraction, which often overlooks fine-grained details in the scene. As a result, these models struggle to detect events that occupy only a small spatial and temporal region of the video, such as ball positions or yellow/red card events. To address this issue, the authors proposed the Unifying Global and Local (UGL) Module, designed to capture both global environmental context and local features from task-relevant scene entities. For extracting global features, UGL adopts the same approach as the E2ESpot model [15], utilizing RegNet-Y with GSM. To capture local features, the model integrates GLIP [57], a vision-language model that processes a single image and a textual prompt describing the target objects to detect (e.g., ”ball,” ”referee,” ”card”). This represents the first use of vision-language models in PES, enabling the model to better focus on semantically relevant entities within the scene. By effectively combining global and local information, the UGL module achieved SOTA performance at the time of publication, demonstrating the efficacy of integrating vision-language models and hybrid feature modeling for fine-grained event detection in sports videos.
Recognizing the need for high temporal resolution in PES, A. Xarles et al. later introduced the Temporal Discriminability Enhancer Encoder-Decoder (T-DEED) [43]. T-DEED addresses the challenge of distinguishing between visually similar frames that occur in close succession by preserving fine-grained temporal resolution across multiple scales. It incorporates Scalable-Granularity Perception (SGP) layers [58] to boost intra-sequence feature discriminability.
@@ 3.2 Action Spotting & Precise Event Spotting
Despite these advances, one major challenge persists: the reliance on large volumes of annotated data. To address this, S. Giancola et al. proposed the first active learning framework for AS [59]. Their pipeline begins with a model trained on a small labeled set, which is then used to generate predictions on unlabeled data. These predictions are ranked using uncertainty and entropy measures to identify the most informative samples. Selected clips are manually annotated and added back into the training loop. This iterative process significantly reduces the required labeled data while maintaining performance—achieving results comparable to state-of-the-art methods using only one-third of the dataset.
@@
Most recently, J. Denize et al. proposed COMEDIAN, the first AS-specific framework that combines self-supervised learning (SSL) with knowledge distillation (KD) for initializing spatiotemporal transformers [44]. COMEDIAN features two Transformers: a spatial transformer that learns short-range context within video segments, and a temporal transformer that integrates longer-range context. The spatial transformer is trained using Momentum Contrast (MoCo) [60], a contrastive learning method that promotes feature similarity across related clips. Both components are further enhanced using a Soft Contrastive loss (SCE) [61], which distills knowledge from a bank of rich spatiotemporal features generated by a pre-trained video model. After pretraining, the entire framework is fine-tuned in a supervised manner for AS, achieving state-of-the-art results on the SoccerNet-v2 benchmark.
3.3 Evaluation Metrics
@@
3.3.1 Mean Average Precision (mAP)
For TAL and AS tasks, the most widely used metric is Mean Average Precision (mAP), which measures how accurately models localize and classify actions. Given a ranked list of predicted intervals with associated confidence scores, precision and recall are defined as:
@@ 3.3.1 Mean Average Precision (mAP)
3.3.2 Average Recall at Average Number of Proposals (AR@AN)
@@
3.3.3 Area Under the AR-AN Curve (AUC)
@@
3.3.4 Tolerance Windows and AP Calculation in AS and PES
@@
3.4 Datasets
@@
In this section, we review publicly available sports-related datasets commonly used for event detection. It is also important to note that although some datasets are originally annotated with temporal intervals, they can often be adapted for AS and PES tasks by treating the start or end frame of each interval as a discrete event label.
@@ 4.1 Generalization Across Diverse Sports
Most existing AS and PES models are developed and validated primarily within individual sports contexts, such as soccer (e.g., SoccerNet). These specialized approaches perform well within their target sports but often exhibit limited capability for generalization, as they rely on domain-specific characteristics like camera angles, event definitions, and player interactions.
Only a limited number of approaches have systematically evaluated their effectiveness across multiple sports datasets (see Table 2), including tennis, figure skating, diving, and gymnastics. This gap highlights a significant limitation: current models struggle to generalize to sports with different visual characteristics, dynamics, and temporal structures. Future research must prioritize:
  • Development of universal AS and PES models leveraging cross-domain transfer learning, possibly employing large-scale pre-trained vision-language models (e.g., CLIP [75]), to enhance cross-sport adaptability.
  • Rigorous benchmarking across diverse sports to validate generalizability, robustness, and practical utility in varied real-world scenarios.
  • The creation and release of new annotated datasets spanning a wider range of sports—particularly underrepresented domains such as volleyball, hockey, fencing, and athletics—to support broader generalization and fair evaluation.
4.2 Advancements in Unsupervised and Low-Supervision Methods
@@ 5 Conclusion
Our analysis highlights the urgent need for the development of generalized frameworks capable of robust performance across multiple sports disciplines, potentially leveraging transfer learning and cross-domain evaluation strategies. Moreover, advancing unsupervised and low-supervision approaches, such as self-supervised learning and active learning, will be critical to reducing annotation costs and enabling scalable deployment. In parallel, future research should explore sophisticated multimodal fusion techniques that effectively integrate visual, audio, and textual modalities to enhance temporal precision, robustness, and event understanding in complex real-world environments.

### v2

@@ 3.2.2 Encoder-Based Methods
Table 2: Summary of methods for Action Spotting (AS) and Precise Event Spotting (PES). Performance is grouped under the Test and Challenge sets. All results are reported on the SoccerNet Action Spotting dataset, where italicized entries indicate results from SoccerNet-v1 [20], and plain text entries are from SoccerNet-v2 [57]. Bold numbers indicate the highest scores in each column. A ✓in the last column means the method was evaluated on datasets beyond SoccerNet. C.D. Eval means cross-dataset evaluation.
Method Year Category Parameter Size Test Set Challenge Set C.D Eval.
@@
Giancola et al. [20] 2018 Pooling-Based 31.37 30.74
Rongved et al. [36] 2020 End-to-End Temporal Learning 56.3
Vats et al. [48] 2020 Temporal Encoder-Based 60.1
CALF [46] 2020 Pooling-Based 41.61 42.22
Vanderplaetsen et al. [58] 2020 Multi-Modal Fusion 39.90
NetVLAD++ [38] 2021 Pooling-Based 53.40 52.54
RMS-Net [49] 2021 End-to-End Temporal Learning 63.49 60.92
Zhou [39] 2021 Pooling-Based 47.05 74.77 49.56 74.84
E2E-Spot [47] 2022 End-to-End Temporal Learning 4.5M 66.73 73.62
Shi et al. [56] 2022 Temporal Encoder-Based 55.20
STE [53] 2022 Temporal Encoder-Based 2.3M 58.29 71.58 58.71 70.49
SpotFormer [50] 2022 Temporal Encoder-Based 60.90 81.50
Soares et al. [54] 2022 Temporal Encoder-Based 8.9M 65.07 78.59 68.33 78.06
Zhu et al. [45] 2022 Pooling-Based 52.04 60.86
ASTRA [59] 2023 Multi-Modal Fusion 70.10 79.21
T-DEED [60] 2024 End-to-End Temporal Learning 16.4M
COMEDIAN [61] 2024 Data-Efficient Learning 29.1M 73.10 68.38 73.98
Tran et al. [62] 2024 End-to-End Temporal Learning 62.49 73.98 69.38 76.15
3.2.3 Multi-Modal Based Methods
Multi-Modal Based Methods extend visual modeling by incorporating complementary modalities, such as audio. Acoustic cues—like whistles, ball strikes, or crowd reactions—often correlate with event boundaries and offer additional context.
Vanderplaetsen et al. [58] evaluated several fusion strategies and found that late fusion, where audio and visual features are combined just before classification, yielded the best performance on SoccerNet.
Expanding on this, Xarles et al. [59] introduced ASTRA, a Transformer-based encoder-decoder model that jointly processes audio and visual embeddings using learnable cross-modal queries. ASTRA achieved strong performance, showcasing the power of cross-modal attention in sports event detection.
3.2.4 Frame-Discriminative Models
Frame-Discriminative Models aim to enhance the spatiotemporal representational capacity of networks by modifying the base architecture or training strategy, often incorporating frame-specific mechanisms to improve temporal discriminability and support precise event spotting. Unlike pooling and encoder based methods, which rely on models designed for unrelated tasks, these approaches allow the network to jointly learn low-level visual features and high-level temporal semantics tailored to the spotting task, resulting in more discriminative and temporally precise predictions.
Hong et al. [47] proposed E2E-Spot, an end-to-end trainable model built on RegNet-Y for visual encoding. To capture fine-grained temporal dynamics, they integrated Gate Shift Modules (GSM) [63], which explicitly model temporal shifts at the frame level. A bidirectional GRU [64] is then employed to process the sequential features and generate per-frame event probabilities, enabling frame-accurate localization suitable for PES.
To overcome the limitations of global-only modeling in E2E-Spot, Tran et al. [62] introduced the Unifying Global and Local (UGL) module. While maintaining the RegNet-Y + GSM backbone for global spatiotemporal encoding, UGL incorporates GLIP [65], a vision-language model, for fine-grained local semantic reasoning (e.g., detecting contextual cues like “referee” or “ball”). This fusion of global temporal encoding with localized semantic awareness enhances the model’s ability to detect subtle or ambiguous events such as fouls or off-screen actions, achieving state-of-the-art results on SoccerNet.
To further improve temporal resolution and disambiguate closely spaced events, Xarles et al. proposed T-DEED [60], a Transformer-based encoder-decoder architecture specifically designed for PES. It introduces Scalable-Granularity Perception (SGP) layers [66] and Gate Shift Fusion (GSF) modules [67], enabling the model to retain temporal discriminability across tightly clustered frames. This design proves especially effective in scenarios with rapid, visually similar events—common in fast-paced sports settings requiring high precision.
3.2.5 Other Models
@@
Despite significant progress, a key bottleneck in AS and PES remains the reliance on large-scale annotated datasets. To address this, Giancola et al. [68] proposed the first active learning framework for action spotting. Their approach begins with a model trained on a small labeled subset, which is then used to generate predictions on unlabeled data. These predictions are ranked by uncertainty and entropy-based heuristics to identify the most informative samples. Manually annotated clips are iteratively added back into the training pool, reducing annotation effort while maintaining competitive performance—achieving results comparable to state-of-the-art models with only one-third of the labeled data.
@@
More recently, Denize et al. introduced COMEDIAN [61], the first AS-specific framework to combine self-supervised learning (SSL) and knowledge distillation (KD) for pretraining spatiotemporal Transformers. The architecture includes a spatial transformer that captures short-range local context and a temporal transformer that models long-range dynamics. The spatial component is trained using Momentum Contrast (MoCo) [69], encouraging consistent representations across temporally adjacent clips. Additionally, a Soft Contrastive Loss (SCE) [70] is used to distill knowledge from a rich feature bank generated by a pre-trained video model. After the SSL phase, the network is fine-tuned with supervision on the AS task, yielding state-of-the-art performance on SoccerNet-v2 while reducing reliance on labeled data.
While existing methods span a range of architectural strategies, each has trade-offs. Pooling-based methods offer computational simplicity by aggregating features over time, though this comes at the cost of temporal precision. Temporal encoders improve contextual modeling but often rely on generic, non-specialized features. Multi-modal methods enhance detection with audio or text cues but depend on modality alignment, while data-efficient approaches reduce labeling costs yet remain limited by their supervision strategies.
A common limitation is the reliance on backbones designed for unrelated tasks like image classification or coarse action recognition, which lack frame-level discriminability and lead to fragmented pipelines.
This highlights the growing need for Frame-Discriminative frameworks that jointly learn spatial and temporal features from raw inputs. Such models promise better precision, adaptability, and generalization—particularly when built with task-specific backbones and unified, modality-aware training objectives.
4 Evaluation Metrics
@@
4.1 Mean Average Precision (mAP)
For TAL tasks, the most widely used metric is Mean Average Precision (mAP), which measures how accurately models localize and classify actions. Given a ranked list of predicted intervals with associated confidence scores, precision and recall are defined as:
@@ 4.1 Mean Average Precision (mAP)
4.2 Average Recall at Average Number of Proposals (AR@AN)
@@
4.3 Area Under the AR-AN Curve (AUC)
@@
4.4 Tolerance Windows and AP Calculation in AS and PES
@@
A key limitation of current AS/PES evaluation is its sensitivity to low confidence thresholds and multi-label predictions. Allowing multiple predictions per frame (e.g., with a 0.1 threshold) increases the chance of hitting the tolerance window and boosting AP, even if other incorrect predictions don’t contribute false positives—resulting in inflated mAP scores. This favors over-predictive models and may obscure true discriminative performance, especially in datasets with sparse or temporally isolated events.
To address this, future benchmarks should adopt stricter per-frame constraints—such as top-1 filtering, calibrated thresholds, or penalties for multi-class outputs—to ensure scores better reflect practical spotting accuracy.
While AS and PES share similar AP computation, the tolerance window and prediction filtering choices play a critical role in shaping fairness and interpretability.
5 Datasets
@@
In this section, we review publicly available sports-related datasets commonly used for event detection. It is also important to note that although some datasets are originally annotated with temporal intervals, they can often be adapted for AS and PES tasks by treating the start or end frame of each interval as a discrete event label. An overview of these datasets is presented in Table 3.
@@ 5.0.4 Other Sports
One core limitation of current datasets is that most are recorded from professionally broadcasted footage with high-quality cameras, ideal angles, and stable viewpoints. While valuable for research, this setting does not reflect the reality faced by many sports analysts and practitioners, especially at non-elite or amateur levels. In many cases, analysts must rely on self-recorded videos with limited camera coverage, occlusions, and inconsistent viewpoints. As a result, models trained solely on broadcast footage may struggle to generalize to practical deployment scenarios where recording conditions are far less controlled.
@@ 6.1 Generalization Across Diverse Sports
While many AS and PES models achieve strong results within individual sports—particularly soccer, as showcased by SoccerNet—they often rely heavily on domain-specific visual and contextual cues such as camera angles, common action semantics, and gameplay structure. This reliance limits their transferability to sports with different visual dynamics, motion patterns, and temporal scales.
A core limitation in current methods is the dependence on backbone architectures originally developed for image classification or coarse-grained action recognition. These models typically process video in fixed-length chunks and apply spatial-temporal aggregation, which diminishes subtle frame-level distinctions that are essential for spotting tasks. In contrast, AS and PES require temporally fine-grained representations that can detect minimal variations between adjacent frames—such as a foot making contact with a ball, or a player entering the field.
To improve generalization and robustness, future work should prioritize the development of frame-level feature extractors tailored specifically to the demands of sports video spotting. Promising directions include:
  • Designing encoders that preserve local temporal granularity, using lightweight 1D CNNs, temporal contrastive learning, or frame-attentive modules to enhance discriminative capacity.
  • Leveraging multimodal pretraining (e.g., CLIP [85]) to align visual, textual, and audio cues into semantically rich frame embeddings suitable for cross-domain transfer.
  • Exploring adaptive frame sampling strategies that focus representational capacity on moments of high temporal importance, improving both efficiency and localization accuracy.
By enhancing frame-wise representation learning, future AS and PES models will be better equipped to operate across a broader range of sports scenarios, improving their temporal precision and reducing reliance on domain-specific heuristics.
6.2 Unsupervised and Low-Supervision Methods
@@ 6.3 Enhanced Multimodal Fusion Approaches
6.4 Real-World Applications: Gaps in Datasets and Evaluation Protocols
Despite impressive progress in AS and PES research, a substantial gap remains between academic benchmarks and real-world deployment scenarios. Most existing datasets are curated from professionally broadcasted footage, captured using high-quality cameras, ideal lighting, and fixed angles. While this enables consistent annotations and controlled evaluations, it fails to reflect the challenges encountered by sports analysts, coaches, and practitioners operating outside elite or televised settings. In many cases, especially at the amateur or semi-professional level, analysts rely on self-recorded videos using handheld devices or static single-angle setups, often under suboptimal conditions. Models trained exclusively on curated datasets may struggle to generalize to such noisy, diverse input sources.
In addition to data quality limitations, current evaluation metrics also fall short of representing practical constraints. Many AS and PES benchmarks compute mean mAP using low confidence thresholds (e.g., 0.1), allowing multiple class predictions per frame. While this boosts recall and improves mAP by capturing more true positives, it introduces an inflated performance signal—particularly problematic in domains such as racket sports, where only one event (e.g., hit, bounce) occurs at a time. In such cases, multi-label predictions per frame are unrealistic and may even degrade performance in downstream applications like rule enforcement or tactical feedback.
To address these shortcomings, future research should aim to:
  • Develop and evaluate models on datasets recorded in unconstrained, real-world environments to ensure robustness beyond broadcast-quality footage.
  • Introduce evaluation protocols that penalize over-prediction and reward frame-level discriminability—e.g., top-1 class selection per frame, or class-wise calibrated thresholds that reflect deployment precision requirements.
  • Encourage the release of datasets collected across a wider range of venues, competition levels, and camera setups to bridge the domain gap between research and practice.
Closing these gaps between academic settings and real-world use cases is critical for building sports video analysis systems that are not only accurate in theory but also reliable, efficient, and trustworthy in practice.
@@ 7 Conclusion
Our analysis underscores the pressing need for the development of frame-level, generalizable frameworks that can robustly operate across diverse sports domains. Achieving this will likely require advances in task-specific backbone design, as well as the cross-domain evaluation strategies to ensure adaptability and performance consistency in varied real-world scenarios. Moreover, advancing unsupervised and low-supervision approaches, such as self-supervised learning and active learning, will be critical to reducing annotation costs and enabling scalable deployment. In parallel, future research should explore sophisticated multimodal fusion techniques that effectively integrate visual, audio, and textual modalities to enhance temporal precision, robustness, and event understanding in complex real-world environments.